The curre Earth, called plate tectonics, tells us that the continents are in constant motion. Assume that the North American continent can be represented by a slab of rock 4700 km on a side and 28 km deep and that the rock has an average mass density of 2940 kg / (m ^ 3) The continent is moving at the rate of about 4.2 cm / y * e * a * r . What is the mass of continent ? Answer in units of kg
What is the kinetic energy of the continent
a jogger (75kg) has the same kinetic energy what would his speed be?
1
Expert's answer
2020-12-17T07:09:32-0500

a)     "Mass= average density(\u2113)* slab rock (a^2) * depth (h)\n\nMass= (\u2113)(a^2) (h)" "Mass= average density(\u2113)* slab rock (a^2) * depth (h")

"Mass= (\u2113)(a^2) (h)" Mass= (ℓ)(a^2) (h)

Mass of the continent= (2940) (4700000) ^2 (28000)

                                    ="1.818 * 10^21 KG"



b)     "Kinetic energy= \u00bd MV^2\n\n =\u00bd * (1.818 * 10^21) (4 * 10^-2) ^2\n\n =1.4544 * 10^18 J" "Kinetic energy= \u00bd MV^2" "Kinetic energy= \u00bd MV^2"

                         ="\u00bd * (1.818 * 10^21) (4 * 10^-2) ^2"

= "\u00bd * (1.818 * 10^21) (4 * 10^-2) ^2"

                           ="1.4544 * 10^18 J"  



c)     Speed of the jogger having the same kinetic energy as that of the continent

Hence K. E= 1.4544 * 10^18 J

Mass= "75kg"

"From the K. E=\u00bd MV^2"

"V= \u221a2 * 1.4544 * 10^18 J\/75"

"V=7.638 *10^8" "m\/s"